OpenStudy (anonymous):

help? Sandra uses her office fax machine to send fax at the rate of $0.10 per page. She decides to rent a fax machine for $80 a year. The cost of sending a fax using the rented machine is $0.06 per page.

OpenStudy (solomonzelman):

what is the question?

OpenStudy (anonymous):

i need help with part a @SolomonZelman Part A: Write an inequality that can be used to calculate the number of pages that Sandra should fax in a year so that the amount she pays for the rented machine is less than the office machine. Define the variable used.

OpenStudy (solomonzelman):

Lets work this together, OK?

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

1) let x represent number of pages 2) each page cost .10+.06 3) additional $80 dollars per year I don't exactly understand the question though....

OpenStudy (solomonzelman):

"the amount she pays for the rented machine is less than the office machine" what do you mean by less than office machine?

OpenStudy (anonymous):

she makes 80.16 per year? @SolomonZelman

OpenStudy (solomonzelman):

Oh, my bad I finally understood the question!

OpenStudy (solomonzelman):

Look the w/o the machine she pays .10 per page, with it she pays .6

OpenStudy (solomonzelman):

@lonelyflowers so far so good?

OpenStudy (anonymous):

I'm not getting it :/

OpenStudy (anonymous):

can you tell me how i could calculate this?

OpenStudy (solomonzelman):

Do you get why I said that w/o the machine she pays .10 per page, with it she pays .6?

OpenStudy (anonymous):

no, could you tell me why? im sorry im really bad in math this is my first yr in highschool

OpenStudy (solomonzelman):

THIS IS YOUR PROBLEM. help? Sandra uses her office fax machine to send fax at the rate of $0.10 per page. She decides to rent a fax machine for $80 a year. The cost of sending a fax using the rented machine is $0.06 per page. look at: "to send fax at the rate of $0.10 per page" "cost of sending a fax using the rented machine is $0.06 per page." see?

OpenStudy (anonymous):

yes.

OpenStudy (solomonzelman):

OK, the machine is $80 dollars.

OpenStudy (solomonzelman):

Office machine is .10 per page & Rental machine is .6 per page

OpenStudy (anonymous):

ok. so how can i write an inequality for this?

OpenStudy (solomonzelman):

The difference per page .04 in other words, every page she saves $.04 (using rental machine)

OpenStudy (solomonzelman):

so let x represent the number of pages and 80 dollars the amount that is supposed to be paid off.

OpenStudy (solomonzelman):

.04x>80

OpenStudy (solomonzelman):

@lonelyflowers with me? can you do it now?

OpenStudy (anonymous):

Every time i calculate it, I get 80.16 this is really difficult for me im not getting it at all :/

OpenStudy (solomonzelman):

Do you see how I got the equation above?

OpenStudy (anonymous):

.04*80=32?

OpenStudy (solomonzelman):

Can you tell me whether you understood why is the equation '.04x>80' ?

OpenStudy (anonymous):

are you talking about greater than?

OpenStudy (solomonzelman):

do you see my color now?)

OpenStudy (solomonzelman):

About everything, the .04x>80 do you understand that?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

So once you got the equation- \[.04x>80\]\[.04x \div .04> 80 \div .04\] same as (b/c 25 is a multiplicative inverse of .04)\[.04x \times 25=80 \times 25\]\[x>80 \times 25\]\[x>2000\]\[(•‿•) \]

OpenStudy (anonymous):

so i put this for part a?

OpenStudy (solomonzelman):

Basically. If you get what I did. I'll ask a couple questions to make sure you understood, okay?

OpenStudy (anonymous):

ok.

OpenStudy (solomonzelman):

where did I get the number '0.04' from? where do I get .04x from? why do I set it to equal more than 80?

OpenStudy (anonymous):

every page she saves is 0.04 x represents .04x and you set equal more than 80 because she rents the machine for 80 a year?

OpenStudy (solomonzelman):

Yup! and the reason that .04x is more than 80 rather than less than 80, is b/c the q was asking for "...rented machine is less than ..." right?

OpenStudy (anonymous):

yes.

OpenStudy (solomonzelman):

\[(•‿•) \]

OpenStudy (solomonzelman):

Part A, that's it?

OpenStudy (anonymous):

ok so what do i put for part a now that we solved this? and theres a part b and then im finished.

OpenStudy (anonymous):

well the equation you wrote do i put that for part a ?

OpenStudy (solomonzelman):

for the final answer, x>2000

OpenStudy (solomonzelman):

okay?

OpenStudy (solomonzelman):

What is the Part b?

OpenStudy (anonymous):

ok so x>2000 for part a this is part b Part B: How many pages should Sandra fax in a year to justify renting the fax machine? Show your work.

OpenStudy (solomonzelman):

in other words, after how many pages will the cost of renting the achine be paid off. (in Economic terms).

OpenStudy (solomonzelman):

I misread the question, I already gave you the part b, it is 2000 or more. for part a....

OpenStudy (solomonzelman):

04.x>80 that is the equation that can be used to calculate... as I showed the steps above.

OpenStudy (solomonzelman):

So Part A is .04x>80 & Part B is (at least) 2000

OpenStudy (anonymous):

Ok. thanks!

OpenStudy (anonymous):

i can put this for part A? .04x>80 .04x÷.04>80÷.04 same as (b/c 25 is a multiplicative inverse of .04) .04x×25=80×25 x>80×25 x>2000

OpenStudy (anonymous):

okay nevermind i get it now, thanks :)

OpenStudy (solomonzelman):

Ok, but have you understood, b/c that's why we are here.